# Free Majorana Fermion Meets Gauged Ising Conformal Field Theory on the Fuzzy Sphere

### v2

@@ 6 Gauged Ising CFT
We then analyse the energy spectrum of the Wilson line endpoint. As the gauge charge is locked with $N_{mf}$, we cannot access the bulk spectrum and Wilson line endpoint spectrum at the same time and use the wavefunction overlap scheme [83, 84, 30] to extract the scaling dimension of endpoint operators. Instead, we fix the lowest endpoint operator $\sigma$ at $\Delta=0.518$, and use the second lowest operator $\partial^{\mu}\sigma$ at $\Delta=1.518$ as the calibrator. The lowest operators up to $\Delta<5$ and $l\leq 3$ can be identified with $\mathbb{Z}_{2}$-odd operators in the Ising CFTs, viz. $\sigma$ with $\Delta=0.518$ and its descendants, and $\sigma^{\mu\nu}$ with $\Delta=4.180$ (Figure 9b). Like for the gauge-neutral sector, we fix $t=0.1$ and optimise a cost function consisting $\sigma$, $\partial^{\mu}\sigma$, $\partial^{\mu}\partial^{\nu}\sigma$, $\Box\sigma$, $\Box\partial^{\mu}\sigma$ and $\sigma^{\mu\nu}$ for each system size to find the optimal $\mu$ (Table 1, right panel). The spectra of identified conformal operators for each system size are plotted in Figure 8b. They agree considerably well with the odd sector of Ising CFT.
@@ 7 Discussion
In this paper, we have studied the free Majorana fermion and the gauged Ising CFT on the fuzzy sphere. We consider the setup of one bosonic and one fermionic lowest Landau level with a $1/2$ angular momentum mismatch, and allow the pair conversion between them. By tuning a relative chemical potential, between a fermionic integer quantum Hall (fIQH) phase at high fermion density and a bosonic Pfaffian (bPf) phase ($\mathrm{SU}(2)_{2}$ Chern-Simons theory) at high boson density, we find an intermediate phase described by an $f$-wave chiral topological superconductor (TSC). The three phases are separated by two continuous phase transition. Supported by numerical and theoretical evidence, we show that the fIQH-TSC transition is described by a free Majorana fermion CFT. Through state-operator correspondence, we find that the energy spectrum agrees well with the operator content of free Majorana fermion up to $\Delta\leq 4$. The local fermion operator can be realised by the combination $f^{\dagger}(\mathbf{r})b(\mathbf{r})$ on the sphere, whose two-point correlation function agrees with the conformal correlator. On the other hand, the TSC-bPf transition is described by a gauged Ising (Ising) CFT. Its spectrum agrees with the $\mathbb{Z}_{2}$-even sector of the Ising CFT. The transition is only consistent with even $N_{mf}$. We find that odd $N_{mf}$ is equivalent to the insertion of a $\mathbb{Z}_{2}$ gauge charge, and the spectrum corresponds to the Wilson line endpoint operators of the gauged Ising CFT, i. e. the $\mathbb{Z}_{2}$-odd sector of the Ising CFT.
This is the first time that a CFT with critical local fermion has been realised on the fuzzy sphere. Other more interesting interacting fermionic CFTs could be studied in the future using similar setup. The simplest target would be the super-Ising CFT, which is a single Majorana fermion coupled to a real Yukawa field $\sigma$. The $\mathbb{Z}_{2}$ symmetry of $\sigma$ is locked with parity due to the Yukawa coupling $\sigma\bar{\chi}\chi$. This theory has enhanced $\mathscr{N}=1$ supersymmetry, where $\sigma$, $\chi$ and $\epsilon=\sigma^{2}$ combines into a superconformal multiplet. Our current setup does not have a manifest parity symmetry, so one has to tune at least three relevant operators, viz. $\sigma$, $\epsilon$, and $\sigma^{\prime}$. Finding a way to implement the parity for Majorana fermion would greatly help targeting this CFT. Another noticeable direction is to generalise this phase diagram as $\mathrm{SU}(2)_{1}$ coupled to an adjoint fermion to other adjoint QCDs [65], which may describe more exotic phase transitions.
In addition, Chern-Simons-matter theories with fermionic content is also an interesting target for fuzzy sphere. For example, the confinement transition of $\nu=1/3$ Laughlin state which describes its transition with a trivially gapped, has $2\pi$-monopole with spin-$3/2$ [85] which may be enhanced to the supercurrent in a $\mathscr{N}=2$ supersymmetry, but the dual supersymmetric theory is expected to be gapped [86]. Another series of examples are the QED with Chern-Simons coupling [20], which describe transitions between Jain states [87]. For example, the $N_{f}=3$ QED describes the transition between the states with fillings $\nu=1/3$ and $\nu=2/3$, and has $4\pi$-monopoles with spin-$1/2$ and $3/2$. Fuzzy sphere is a promising platform to study these theories, as the realisation of the proximate topological orders are natural with the fractional quantum Hall setup, e. g., the Laughlin states [88] with filling $\nu=1/k$ realise the Abelian topological orders captured by Chern-Simons theories $\mathrm{U}(1)_{-k}$, the Jain sequence [87] with $\nu=p/(mp+1)$ realises the Abelian topological orders with composite fermion descriptions, and the Read-Rezayi sequence with $\nu=p/(mp+2)$ [89], as a natural extension of the Moore-Read states [54], realises in general a series of non-Abelian topological orders.
@@
Our work also represents a step towards studying renormalisable Lagrangians of quantum field theories on the fuzzy sphere. Most of the fuzzy sphere works so far are limited to realising universalities by constructing Hamiltonian with correct symmetry and anomaly and letting it flow to the desired conformal fixed point in the IR. In the meantime, realising renormalisable Lagrangians within the fuzzy sphere framework will be the cornerstone towards QFT questions like field theory duality. A natural starting point is to understand free theories on fuzzy sphere, such as the free fermion CFT studied here as well as free real scalar CFT in Ref. [46]. Our fuzzy sphere Majorana fermion model has several interesting features which deserve further exploration:
@@
    The critical point lies at $\mu_{c1}=0$ at high precision for a wide range of $t$, hinting at some enhanced symmetry that forbids the chemical potential for the free Majorana fermion CFT.
@@
    The Hamiltonian at the critical point $\mu_{c1}=0$ resembles the familiar free Majorana fermion QFT. Specifically, apart from the charge density interaction term $n_{e}^{2}$, the Hamiltonian takes the form $\bar{\chi}i\not{\partial}\chi$, where $\chi$ denotes the Majorana fermion.

### v3

@@ 6 Gauged Ising CFT
We then analyse the energy spectrum of the Wilson-line end-point. As the gauge charge is locked with $N_{mf}$, we cannot access the bulk spectrum and Wilson-line end-point spectrum at the same time and use the wave-function overlap scheme [88, 89, 49] to extract the scaling dimension of end-point operators. Instead, we fix the lowest end-point operator $\sigma$ at $\Delta=0.518$, and use the second lowest operator $\partial^{\mu}\sigma$ at $\Delta=1.518$ as the calibrator. The lowest operators up to $\Delta<5$ and $l\leq 3$ can be identified with $\mathbb{Z}_{2}$-odd operators in the Ising CFTs, viz. $\sigma$ with $\Delta=0.518$ and its descendants, and $\sigma^{\mu\nu}$ with $\Delta=4.180$ (Figure 9b). Like for the gauge-neutral sector, we fix $t=0.1$ and optimise a cost function consisting $\sigma$, $\partial^{\mu}\sigma$, $\partial^{\mu}\partial^{\nu}\sigma$, $\Box\sigma$, $\Box\partial^{\mu}\sigma$ and $\sigma^{\mu\nu}$ for each system size to find the optimal $\mu$ (Table 1, right panel). The spectra of identified conformal operators for each system size are plotted in Figure 8b. They agree considerably well with the odd sector of Ising CFT.
7 Correspondence with QFT Lagrangians
In Ref. [35], an analogy between the fuzzy-sphere models and the field theories have been noted for the free and critical scalars. In this section, we further extend this correspondence to free and interacting Majorana fermions.
For a critical real scalar (Ising), the set-up contains two fermionic flavours $f_{0,1}$ that are respectively even and odd under the $\mathbb{Z}_{2}$ symmetry. The $f_{0}$ flavour can be identified as a ‘vacuum’ or reference flavour — the CFT ground state can be regarded as a perturbation from the reference state where all the $f_{0}$ particles are occupied. As an evidence, the CFT ground state has a substantial overlap with the reference state. As the reference state is a direct product state in the orbital space, the orbital-space entanglement entropy is very small for the CFT ground state. The orbital-space entanglement entropy is measured through a bi-partition between the $m<0$ orbitals and the $m\geq 0$ orbitals. Correspondingly, the $f_{1}$ flavour is identified as an ‘excited state.’ The hopping between the two flavours is identified as the elementary field $\sigma$, and the density of ‘excitations,’ or particles in the $f_{1}$ flavour, is identified as the mass field $\epsilon=\sigma^{2}$ up to a constant
$\sigma$ $\sim n_{x}=f_{1}^{\dagger}f_{0}+f_{0}^{\dagger}f_{1},$ $\epsilon+\text{const.}$ $\sim n_{1}=f_{1}^{\dagger}f_{1}.$ (13)
Here the $\sim$’ denotes a phenomelogical correspondence between a field theory operator and a fuzzy-sphere operator, or between a term in the field-theory Lagrangian and a term in the fuzzy-sphere Hamiltonian. [^12] The correspondence holds more precisely between the field-theory Hamiltonian and fuzzy-sphere Hamiltonian. Since the Lagrangian and Hamiltonian can be easily converted, we pick the Lagrangian as a more familiar representative. The correspondence is illustrated in Figure 10a. All the terms in the Hamiltonian
$H=\int\mathrm{d}^{2}\mathbf{r}\left[n_{e}(\mathbf{r})^{2}+Un_{x}(\mathbf{r})\nabla^{2}n_{x}(\mathbf{r})-\mu_{1}n_{1}(\mathbf{r})\right],$ (14)
besides a electric density interaction of $n_{e}=f_{0}^{\dagger}f_{0}+f_{1}^{\dagger}f_{1}$ that suppresses the charge fluctuation, has a direct intepretation to the field theory
$\mathscr{L}$ $=\frac{1}{2}(\partial^{\mu}\sigma)^{2}-\frac{1}{2}m^{2}\sigma^{2}-\frac{\lambda}{4!}\sigma^{4}.$ (15)
Similar to the lattice models like the transverse-field Ising model, the Ising-type density-density interaction can be intepreted as the $\phi^{4}$-theory at a certain mass $m_{0}$, and the chemical potential fine tunes the mass term
$Un_{x}\nabla^{2}n_{x}$ $\sim\frac{1}{2}(\partial^{\mu}\sigma)^{2}-\frac{1}{2}m_{0}^{2}\sigma^{2}-\frac{\lambda}{4!}\sigma^{4},$ $\mu_{1}n_{1}$ $\sim\frac{1}{2}(m^{2}-m_{0}^{2})\sigma^{2}.$ (16)
Figure 10: The illustration of the correspondence between fuzzy-sphere set-up and the field-theory operators in (a) the critical real scalar (Ising), (b) the free Majorana fermion, and (c) the super-Ising theory.
While the correspondence between field-theory Lagrangian and the fuzzy-sphere model is to some extent vague for the critical scalar, it can be seen much more clearly for the free Majorana fermion. The set-up now contains one fermionic ‘reference’ flavour $f$ and a bosonic ‘excitation’ flavour $b$. The CFT ground state has a substantial overlap $|\langle\text{ref}|0_{\text{CFT}}\rangle|$ with the reference state where all fermions are occupied (Figure 11a)
$|\text{ref}\rangle=\prod_{m=-q}^{q}f_{m}^{\dagger}|\emptyset\rangle$ (17)
and its orbital-space entanglement entropy is small (Figure 11b). The conversion between the two flavours are identified as the elementary spinor field $\chi$, and the density of ‘excitations,’ or particles in the $b$ flavour, is identified as the mass field $\bar{\chi}\chi$ up to a constant (Figure 10b)
$\chi$ $\sim(\eta,\eta^{\dagger})^{\mathrm{T}}=(f^{\dagger}b,b^{\dagger}f)^{\mathrm{T}},$ $\bar{\chi}\chi+\mathrm{const.}$ $\sim n_{b}=b^{\dagger}b$ (18)
The terms in the Hamiltonian can again be intepreted as the kinetic term and the mass term
$\frac{1}{2}\bar{\chi}\,i\not{\partial}\chi$ $\sim t\left(\eta D_{+}\eta+\text{h.\penalty 10000\ c.}\right),$ $\frac{1}{2}m\bar{\chi}\chi\sim\mu n_{b}$ (19)
where $D_{+}=D_{\theta}+iD_{\phi}$ in addition to an electric density interaction.
Notably, the pair conversion term $H_{t}$ bears exactly the same form as the kinetic term in the Hamiltonian of free Majorana fermion upon radial quantisation.
$H_{t}\propto\int\mathrm{d}^{2}\mathbf{r}\,i\epsilon^{\mu\nu\rho}\hat{n}_{\mu}\bar{\chi}_{\alpha}(\mathbf{r})(\gamma_{\nu})^{\alpha}{}_{\beta}\partial_{\rho}\chi^{\beta}(\mathbf{r})=\int\mathrm{d}^{2}\mathbf{r}\,i\bar{\chi}(\mathbf{r})(\hat{\mathbf{n}}\times\gamma\cdot\partial)\chi(\mathbf{r}).$ (20)
As an evidence for this correspondence, the critical point locates at $\mu_{c}=0$ to a high precision where the mass term vanishes.
Figure 11: Properties of the CFT ground state of the free Majorana fermion: (a) The overlap $|\langle\text{ref}|0_{\text{CFT}}\rangle|$ between the CFT ground state and the reference state (17) where all fermions are occupied; (b) the orbital-space entanglement entropy. The data are measured at $t=0.3,\mu_{c1}=0$ and plotted as a function of the system size $N_{mf}$.
We shall show that these ingredients suffice to construct the interacting theories. The simplest one is a Majorana fermion $\chi$ coupled to a real scalar $\sigma$ through a Yukawa interaction [13, 90]
$\mathscr{L}[\chi,\sigma]=\frac{1}{2}\bar{\chi}\,i\not{\partial}\chi+\frac{1}{2}(\partial^{\mu}\sigma)^{2}-\frac{1}{2}m_{\sigma}^{2}\sigma^{2}-\frac{\lambda}{4!}\sigma^{4}-\frac{g}{2}\sigma\bar{\chi}\chi.$ (21)
The Yukawa interaction locks the Ising $\mathbb{Z}_{2}$ symmetry to the space-time parity. As a result, this theory has no global symmetry, and $\sigma$ is odd under space-time parity. When the parity is not manifest, two additional perturbations should also be considered, namely the fermion mass and the polarisation term
$m_{\chi}\bar{\chi}\chi+h\sigma.$
In the infra-red, the Lagrangian flows to a $\mathscr{N}=1$ super-conformal field theory (SCFT) known as the super-Ising theory. We will introduce its property in more detail in Section 8.
The corresponding set-up on the fuzzy sphere contains two fermionic flavours $f_{0},f_{1}$ and a bosonic flavour $b$. We set the bosons and fermions to have a mismatch $q_{f}-q_{b}=1/2$ in the angular momenta. We set the total filling to be $1$, i. e. $N_{e}=N_{mf}$. Similar to the critical real scalar and the free Majorana fermion, we interpret the $f_{0}$ flavour as the reference flavour, $f_{1}$ as the scalar excitation and $b$ as the fermion excitation. Hence, the elementary fields $\chi$ and $\sigma$ are realised as the hoppings from the reference fermion flavour $f_{0}$ to respectively the boson flavour $b$ and the other fermion flavour $f_{1}$ (Figure 10c)
$\eta(\mathbf{r})$ $=b^{\dagger}(\mathbf{r})f_{0}(\mathbf{r}),$ $\chi(\mathbf{r})$ $\sim(\eta(\mathbf{r}),\eta^{\dagger}(\mathbf{r}))^{\mathrm{T}}$ (22a)
$n_{x}(\mathbf{r})$ $=f_{1}^{\dagger}(\mathbf{r})f_{0}(\mathbf{r})+f_{0}^{\dagger}(\mathbf{r})f_{1}(\mathbf{r}),$ $\sigma(\mathbf{r})$ $\sim n_{x}(\mathbf{r})$ (22b)
The fermion and scalar bilinears $\bar{\chi}\chi$ and $\sigma^{2}$ are realised using the density operators
$n_{b}$ $=b^{\dagger}b\sim\bar{\chi}\chi+\text{const.},$ $n_{1}=f_{1}^{\dagger}f_{1}\sim\sigma^{2}+\text{const.}$ (22c)
up to a constant. Using these ingredients, we can translate the terms in the Lagrangian to four-particle interactions in the fuzzy-sphere model. We start with the terms that describe the critical scalar and the free Majorana fermion
$\frac{1}{2}\bar{\chi}\,i\not{\partial}\chi$ $\sim t\left(\eta(\mathbf{r})D_{+}\eta(\mathbf{r})+\text{h.\penalty 10000\ c.}\right)$ (23a)
$\frac{1}{2}(\partial^{\mu}\sigma)^{2}-\frac{1}{2}m_{0}^{2}\sigma^{2}-\frac{\lambda}{4!}\sigma^{4}$ $\sim Un_{x}(\mathbf{r})\nabla^{2}n_{x}(\mathbf{r}).$ (23b)
The Yukawa coupling can be expressed as a contact interaction between $n_{b}$ and $n_{x}$
$\frac{g}{2}\sigma\bar{\chi}\chi\sim gn_{x}(\mathbf{r})n_{b}(\mathbf{r}).$ (23c)
The terms left are the scalar mass, and the two additional terms due to the absence of parity symmetry
$\frac{1}{2}(m_{\sigma}^{2}-m_{0}^{2})\sigma^{2}$ $\sim\mu_{1}n_{1},$ $\frac{1}{2}m_{\chi}\bar{\chi}\chi$ $\sim\mu_{b}n_{b},$ $h\sigma$ $\sim hn_{x}.$ (23d)
Combined with an electric density interaction $n_{e}^{2}$ where $n_{e}=f_{0}^{\dagger}f_{0}+f_{1}^{\dagger}f_{1}+b^{\dagger}b$, the total Hamiltonian reads
$H=\int\mathrm{d}^{2}\mathbf{r}\,\Big[n_{e}^{2}(\mathbf{r})+t\left(\eta(\mathbf{r})D_{+}\eta(\mathbf{r})+\text{h.\penalty 10000\ c.}\right)+Un_{x}(\mathbf{r})\nabla^{2}n_{x}(\mathbf{r})+gn_{x}(\mathbf{r})n_{b}(\mathbf{r})\\ -hn_{x}(\mathbf{r})-\mu_{1}n_{1}(\mathbf{r})-\mu_{b}n_{b}(\mathbf{r})\Big].$ (24)
It contains six parameters $t,U,g,h,\mu_{1}$ and $\mu_{b}$. The latter three $h,\mu_{1}$ and $\mu_{b}$ controls the three relevant perturbations $\sigma,\epsilon=\sigma^{2}$ and $\sigma^{\prime}=\bar{\chi}\chi$, so that we can fix the values of $t,U,g$ and fine tune $h,\mu_{1},\mu_{b}$ to reach the super-Ising fixed point. [^13] In Eq. (3), we write the Hamiltonian in terms of the pseudo-potentials, while it is written in terms of local operators in Eq. (24) for convenience. The difference produces a constant conversion factor.
8 Super-Ising SCFT
We have constructed the model to realise the super-Ising theory on the fuzzy sphere. In this section, we present the numerical evidence for the emergence of super-conformal symmetry and the agreement with conformal bootstrap.
We first discuss the supersymmetry in the field-theory Lagrangian (21). In the infra-red, it flows to a super-conformal field theory (SCFT) with super-charge $\mathscr{N}=1$ known as the super-Ising theory. The supersymmetry is manifest in the Lagrangian at $\lambda=3g^{2}$ and $m_{\sigma}=0$. The action can be organised in terms of a super-field $\Sigma(x,\theta)=\sigma+\bar{\theta}\chi+\frac{1}{2}\bar{\theta}\theta\epsilon$ where $\theta$ and $\bar{\theta}$ are Grassmannian co-ordinates. The $\epsilon$ is an auxilliary field coupled to $\sigma^{2}$. By the equation of motion $\epsilon=\sigma^{2}$. The Poincaré super-charge $Q^{\alpha}$ acts on the fields as
$[Q^{\alpha},\sigma]$ $=\chi^{\alpha},$ $\{Q^{\alpha},\chi^{\beta}\}$ $=-(\gamma^{\mu})^{\alpha\beta}\partial_{\mu}\sigma-\epsilon^{\alpha\beta}\epsilon,$ $[Q^{\alpha},\epsilon]$ $=(\gamma^{\mu})^{\alpha\beta}\partial_{\mu}\chi_{\beta}.$
The super-conformal group is homomorphic to a super-Lie group $\mathrm{OSp}(1|4)$ [91, 92]. More precisely, the (Lorentzian) conformal transformations acting on fermions give the Lie group $\mathrm{Spin}(3,2)\cong\mathrm{Sp}(4,\mathbb{R})$ of real symplectic transformations. Besides that, the super-Lie algebra also contains two Grassmannian generators, i. e. the Poincaré super-charge $Q^{\alpha}$ and the conformal super-charge $S^{\alpha}$, which extend the conformal group to $\mathrm{OSp}(1|4,\mathbb{R})$. These generators organise several conformal multiplets into a super-conformal multiplet in a similar way as how the conformal symmetry orgainses the scaling operators into conformal multiplets [93]. Similar to $P^{\mu}$ and $K^{\mu}$ that act as ladder operators of $D$ increasing and decreasing the scaling dimension $\Delta$ by $1$, $Q^{\alpha}$ and $S^{\alpha}$ act as ‘half-ladder’ operators changing $\Delta$ by $\pm 1/2$. The conformal primaries are classified into super-conformal primairies that are the lowest-weight state of $Q^{\alpha}$, and the super-conformal descendants by acting $Q^{\alpha}$ on the super-primaries. Due to the odd Grassmannian parity of $Q^{\alpha}$, each super-multiplet contains at most 4 conformal multiplets. As an example, the Majorana fermion $\chi^{\alpha}=Q^{\alpha}\sigma$ and $\epsilon=\bar{Q}_{\alpha}Q^{\alpha}\sigma$ are now super-conformal descendants of the super-conformal primary $\sigma$, and their scaling dimensions are not independent $\Delta_{\chi}=\Delta_{\sigma}+1/2$, $\Delta_{\epsilon}=\Delta_{\sigma}+1$. The operator $\bar{Q}Q$ is odd under space-time parity, so $\sigma$ and $\epsilon$ has opposite space-time parity.
The super-Ising SCFT has been studied extensively by perturbative calculation [12, 94] and conformal bootstrap [14, 15, 16]. The two relevant super-primaries $\sigma,\sigma^{\prime}=\bar{\chi}\chi$ has been determined to a high precision. They are parity-odd Lorentz scalars and have scaling dimensions $\Delta_{\sigma}=0.584,\Delta_{\sigma^{\prime}}=2.887$ [16]. Besides that, a spin-$2$ parity-even operator $T^{\prime\mu\nu}$ at $\Delta_{T^{\prime}}\approx 3.28$ has been observed in the extremal spectrum [16].
To verify that our ansatz (24) realises the super-Ising SCFT, we carry out an ED calculation up to $N_{mf}=9$. We fix $t=1.5$, $U=0.25$ and $g=1$ and determine $h$, $\mu_{1}$ and $\mu_{b}$ for each $N_{mf}$ by optimising a cost function defined as the root mean square of the relative deviations of the scaling dimension of all 14 operators with $\Delta<3.1$ and $l\leq 2$ (listed in Figure 12) from the values obtained from the conformal bootstrap [16]. The conformal points for $N_{mf}=9,8,7,6$ are listed in Table 2. The optimal parameters remain relatively stable with the system size. The low-lying spectrum for various $N_{mf}$ are plotted in Figure 12. For each $N_{mf}$, the lowest operators remain close to the expected values. The cost functions remain between $2\%$ and $3\%$.
$N_{mf}$ $h$ $\mu_{1}$ $\mu_{b}$
99 $0.0680$ $0.0809$ $0.0776$
88 $0.0690$ $0.0792$ $0.0757$
77 $0.0698$ $0.0774$ $0.0722$
66 $0.0662$ $0.0793$ $0.0615$
Table 2: The optimal conformal point of the super-Ising SCFT for different system sizes $N_{mf}$. In the calculation, we fix $t=1.5,U=0.25,g=1$.
Figure 12: The low-lying spectrum of the super-Ising SCFT with $\Delta\lesssim 3$ and $l\leq 2$ for different system sizes. The scaling dimensions are calibrated by optimising the cost function. The grey bars denote the bootstrap values. The names of the operators are marked beside the bars. In the calculation, we fix $t=1.5,U=0.25,g=1$ and take $h,\mu_{1},\mu_{b}$ as in Table 2.
Going to slightly higher part of the spectrum (Figure 13), almost all states at $\Delta\lesssim 5,l\leq 3$ can be classified into super-multiplets of $\sigma$, $G$, $\sigma^{\prime}$ and $T^{\prime}$. The scaling dimensions of these operators are close to bootstrap results. We note that states with $\Delta>3.1$ are not included as inputs in the optimisation of the conformal point; nevertheless, they exhibit good consistency with the spectrum of the super-Ising theory. In Appendix E, we give a full spectrum up to $\Delta=6.5$ and identify four more super-multiplets $\epsilon^{\prime\prime},T_{-},[\sigma\chi]_{l=7/2}$ and $\sigma\sigma]_{l=4}$ from the bootstrap results. We have also confirmed a primary $G^{\prime\mu\alpha}$ at $l=3/2$ and $\Delta\approx 5.28$ that is subject to large numerical uncertainty in the conformal bootstrap [16].
Figure 13: The spectrum of super-Ising field theory up to $\Delta\approx 5$ and $l=3$ at $N_{mf}=9$. The symbols marks the measured scaling dimensions and the bars marks the expected values from conformal bootstrap. Most states are identified in the lowest super-conformal multiplets. Each identified state is labelled by the measured and expected (in parentheses in subscript) scaling dimensions. The brown circles denote higher super-conformal states and three unidentified states. Higher spectrum is given in Appendix E. In the calculation, we fix $t=1.5,U=0.25,g=1$ and take $h,\mu_{1},\mu_{b}$ as in Table 2.
Two key pieces of evidence for the emergent super-conformal symmetry are the supersymmetry Noether current and the super-conformal multiplet structure. For $\mathscr{N}=1$ SCFTs, apart from the stress tensor, there exists another conserved supersymmetry current $G^{\alpha\mu}$ whose spatial integral gives the super-charge $Q^{\alpha}$. It is fermionic, has Lorentz spin $l_{G}=3/2$, and its scaling dimension saturates the unitarity bound $\Delta_{G}=5/2$.
Figure 14: The super-conformal multiplets of the lowest super-conformal primaries (a) $\sigma$, (b) $G$, (c) $\sigma^{\prime}$, and (d) $T^{\prime}$ at $N_{mf}=9$ and up to $\Delta=6$ and $l=4$. The symbols marks the measured scaling dimensions and the bars marks the expected values from conformal bootstrap. The states in each super-conformal multiplets are organised into conformal multiplets. The degenerate states are strongly mixed and are assigned arbitrarily. In the calculation, we fix $t=1.5,U=0.25,g=1$ and take $h,\mu_{1},\mu_{b}$ as in Table 2.
The super-conformal symmetry organises the scaling operators into super-conformal multiplets. Acting $Q^{\alpha}$ on a super-conformal primary gives its super-conformal descendants. The action alters the fermion parity, changes the Lorentz spin by $1/2$ and increases the scaling dimension by $1/2$. For each low-lying super-primary, we can find numerically all of its super-conformal descendants and their conformal descendants up to $\Delta\lesssim 6$ and $l\leq 4$ (Figure 14). Their scaling dimensions are at half-integer spacings as expected. In the following, we explain the structure of the super-multiplets:
For a general super-conformal primary $\Phi$ with Lorentz spin 0, its super-multiplet contains three conformal primaries, viz. the bosonic $\Phi$, fermionic $Q^{\alpha}\Phi$ with spin-$1/2$ and $\Delta_{\Phi}+1/2$, and bosonic $\bar{Q}_{\alpha}Q^{\alpha}\Phi$ with spin-0 and $\Delta_{\Phi}+1$. As an example, the primaries $\sigma$, $\chi=Q\sigma$ and $\epsilon=\bar{Q}Q\sigma$ forms a super-conformal multiplet and their scaling dimensions are locked $\Delta_{\chi}=\Delta_{\sigma}+1/2,\Delta_{\epsilon}=\Delta_{\sigma}+1$ (Figure 14a). For a spin-$l$ primary $\Phi^{\alpha_{1}\dots\alpha_{2l}}$ (where each Lorentz index is expressed in terms of two spinor indices), e. g. the spin-$2$ $T^{\prime}_{\mu\nu}$, its super-multiplet contains $\Phi$, $Q\Phi=Q^{\alpha_{0}}\Phi^{\alpha_{1}\dots\alpha_{2l}}$, $\bar{Q}\Phi=\bar{Q}_{\alpha_{1}}\Phi^{\alpha_{1}\dots\alpha_{2l}}$ with spin-$(l\pm 1/2)$ and $\Delta_{\Phi}+1/2$, and $\bar{Q}Q\Phi$ with spin-$l$ and $\Delta_{\Phi}+1$ (Figure 14d). States from different conformal multiplets can be degenerate protected by the super-conformal symmetry. E. g., we have observed a nearly degeneracy of the conformal primary $\bar{Q}_{\alpha}Q^{\alpha}T^{\prime\mu\nu}$ and the conformal descendant $\epsilon^{\mu\rho\sigma}\partial_{\rho}T^{\prime\nu}{}_{\sigma}$ at $\Delta\approx 4.28$ and $l=2$, although subject to level repulsion [34]. An exception to this structure is the shortening of the super-current super-multiplet due to the conservation $\bar{Q}_{\alpha}G^{\alpha\mu}=0$. It only contains the supersymmetry current $G^{\alpha\mu}$ and the stress tensor $T^{\mu\nu}=Q^{\alpha}(\gamma^{\mu})_{\alpha\beta}G^{\nu\beta}$ as its super-descendant (Figure 14b).
@@ 9 Discussion
In this paper, we have studied the free Majorana fermion, the gauged Ising CFT and the super-Ising SCFT on the fuzzy sphere. For the free Majorana fermion, we consider a set-up consisting of one bosonic and one fermionic lowest Landau level with a $1/2$ angular momentum mismatch, and allow the pair conversion between them. This microscopic realisation of the spin-statistics constraint expected in the IR is an interesting subject for future work. By tuning a relative chemical potential between a fermionic integer quantum Hall (fIQH) phase at high fermion density and a bosonic Pfaffian (bPf) phase ($\mathrm{SU}(2)_{2}$ Chern-Simons theory) at high boson density, we find an intermediate phase described by a Majorana quantum Hall (MQH) state with Kitaev index $\nu_{K}=3$. The three phases are separated by two continuous phase transition. Supported by numerical and theoretical evidence, we show that the fIQH-MQH transition is described by a free-Majorana-fermion CFT. Using state-operator correspondence, we find that the energy spectrum agrees well with the operator content of free Majorana fermion up to $\Delta\leq 4$. The local fermion operator is realised as the bilinear $f^{\dagger}(\mathbf{r})b(\mathbf{r})$ and h. c. on the sphere, whose two-point correlation function matches the expected conformal correlator. On the other hand, the MQH-bPf transition is described by a gauged Ising (Ising) CFT. Its spectrum agrees with the $\mathbb{Z}_{2}$-even sector of the Ising CFT. The transition is only consistent with even $N_{mf}$. We find that odd $N_{mf}$ is equivalent to the insertion of a $\mathbb{Z}_{2}$ gauge charge, and the spectrum corresponds to the Wilson-line end-point operators of the gauged Ising CFT, i. e. the $\mathbb{Z}_{2}$-odd sector of the Ising CFT.
Furthermore, we have established a correspondence between the fuzzy-sphere model Hamiltonians and the field-theory Lagrangians. In this correspondence, a flavour serves as a reference state and the other serves as an ‘excitation’ flavour. The elementary fields that constitute the field theory are identified as hopping between them. Using this framework, we have translated an interacting fermionic CFT — the super-Ising theory, in which a Majorana fermion $\chi$ is coupled to a real scalar $\sigma$ through a Yukawa interaction — into a fuzzy-sphere model with two fermionic flavours $f_{0,1}$ and one bosonic flavour $b$. The elementary fields are identified with the hopping operators $\sigma=f_{0}^{\dagger}f_{1}+f_{1}^{\dagger}f_{0}$ and $\chi=(b^{\dagger}f_{0},f_{0}^{\dagger}b)^{\mathrm{T}}$. The three tuning parameters admit direct physical intepretations as the relevant operators $\sigma$, $\epsilon=\sigma^{\prime}$ and $\sigma^{\prime}=\bar{\chi}\chi$. We have provided numerical evidence for the emergence of super-conformal symmetry and the agreement with the conformal bootstrap results. The super-conformal multiplet structure with half-integer spacing is clearly observed and most of the low-lying states are identified as SCFT operators.
@@
Our work represents the first fuzzy-sphere realisation of CFTs with critical local fermions and constitutes a step towards studying renormalisable Lagrangians of quantum field theories on the fuzzy sphere. Most previous fuzzy-sphere works were limited to realising universality classes by constructing Hamiltonian with correct symmetry and anomaly and relying on RG flow to reach the desired conformal fixed point in the IR. The free Majorana fermion, the super-Ising theory, and the free and critical real-scalar models in Ref. [35] offer a natural and versatile playground for studying the connection between fuzzy-sphere constructions and quantum field theories. In particular, our fuzzy-sphere Majorana-fermion model has several notable features which deserve further exploration:
@@
    The critical point lies at $\mu_{c1}=0$ with high precision over a wide range of $t$, suggesting the presence of an enhanced symmetry that forbids the chemical-potential perturbation in the free-Majorana-fermion CFT.
@@
    At the critical point $\mu_{c1}=0$, the Hamiltonian closely resembles the familiar free Majorana fermion QFT: Apart from the charge density interaction term $n_{e}^{2}$, it takes the same form as the field theory (20).